how many grams of sodium hydrogen carbonate (MW 84.4) are needed to neutralize stomach acid (pH=2) if we assume all the H+ comes from HCL and the volume of the stomach is 950ml?
Answer #1

HCl + NaHCO3 ----> NaCl + CO2 + H2O

1 mol HCl = 1 mol NaHCO3

pH= 2 , [H3O+] = 10^-2 = 0.01 M

no of mol of H3o+ = HCl = 0.01*0.95 = 0.0095 mol

no of mol of NaHCO3 required = 0.0095 mol

amount of NaHCO3 required = 0.0095*84.4 = 0.802 g